Evaluate each of the following expressions and list the final value of variable x.

1) x = 4-(8%5)*3/(8-6)%3;

2) x = ((9%8)*2)*8/5%(7/3)

3)x= (5-2*9+(3+((9/4)%2*3+6/(8%3))))

4)x =3/6+13%3+2


1
Expert's answer
2021-06-11T17:11:58-0400
Expression 1: 
step 1 = 4-3*3/(8-6)%3;
step 2 = 4-3*3/2 %3
step 3 = 4-9 / 2 %3
step 4 = 4-4 %3
step 5 = 4 - 1
Answer = 3
final value of x = 3


Expression 2:
step 1 = (1*2)*8/5%(7/3)
step 2 = 2 *8/5%(7/3) 
step 3 = 2 *8/ 5%2
step 4 =  16 / 5%2
step 5 = 3 %2
Answer = 1
final value of x = 1


Expression 3:
step 1 = (5-2*9+(3+((9/4)%2*3+6/5)))
step 2 = (5-2*9+(3+( 2%2 *3+6/5)))
step 3 = (5-2*9+(3+( 2%6+6/5)))
step 4 = (5-2*9+(3+( 2 +6/5)))
step 5 = (5-2*9+(3+( 2 +1)))
step 6 = (5-18+(3+( 2 +1)))
step 7 = (5-18+(6))
step 8 = 11 - 18
answer = -7
final value of x = -7




Expression 4:
step 1 = 0 +13%3 + 2 
step 2 = 0 + 1 + 2
step 3 = 1 + 2
answer = 3


final value of x = 3